OpenStudy (anonymous):

I need to find the zeros in this equation. -3x^4 + 27x^2 _ 1200 = 0

OpenStudy (michele_laino):

please change the varible to a new variable, say y=x^2. Your eqution will become as below: \[-3y ^{2}+27 y-1200=0\] which can be solved easily for y

OpenStudy (anonymous):

I'm sorry, I'm not sure I understand. How were you able to get rid of a portion of the exponents just by changing the variable?

OpenStudy (anonymous):

Is this correct? \[-3x^4 +27^2 - 1200 + 1200 = 0 + 1200\]

OpenStudy (anonymous):

\[\sqrt{-3x^4 + 27^2} = \sqrt{1200}\]

OpenStudy (michele_laino):

please you have to apply the standard formula to solve the quadratic equation for y

OpenStudy (anonymous):

That first one was supposed to go second.

OpenStudy (anonymous):

What would be the standard formula again?

OpenStudy (anonymous):

I'm not trying to get an answer out of you I just need some help wrapping my head around this stuff.

OpenStudy (michele_laino):

\[ay ^{2}+by+c=0\] then the standard formula is: \[y=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

Wait, you mean the quadratic formula? Okay, I get it now. But doesn't that help me find only the imaginary zeroes?

OpenStudy (campbell_st):

just looking at the dicriminant its negative so the zeros are complex

OpenStudy (michele_laino):

@ericdegr8 your equation is: \[3y ^{2}-27y+1200=0\] so?

OpenStudy (anonymous):

\[x = \frac{ 27\pm \sqrt{27^{2}-4(3)(-1200)} }{ 2(3) }\]

OpenStudy (anonymous):

\[x = \frac{ 27\pm \sqrt{729 + 14,400} }{ 6 }\]

OpenStudy (anonymous):

Correct?

OpenStudy (campbell_st):

I think if you are using the general quadratic formula on the original equation, you have a couple of errors \[x^2 = \frac{-27 \pm \sqrt{(-27)^2 - 4 \times (-3) \times (-1200)}}{2 \times (-3)}\]

OpenStudy (campbell_st):

opps should be 27 inside the square root.. no the interesting problem is that you will now have to take the square root of the answer to find x

OpenStudy (anonymous):

Okay, so I use the quadratic formula to find x^2, then I find the square root of that?

OpenStudy (anonymous):

But what about the real solutions? I thought the quadratic formula only finds complex solutions? Don't you have to factor normally to find the real solutions?

OpenStudy (anonymous):

Or are there any real solutions?

OpenStudy (campbell_st):

the quadratic formula will find all solutions... but checking the discriminant when I 1st saw the question, it was negative so the solutions for x^2 will be complex... the discriminant is \[\Delta = b^2 - 4ac\]

OpenStudy (anonymous):

Okay, so that's the discriminant. This actually clears up a lot of confusion in my math class.

OpenStudy (anonymous):

How about you @magepker728 ? Do you have any input?

OpenStudy (campbell_st):

can I just check the equation is \[-3x^4 + 27x^2 - 1200 = 0\] asyou seem to have a typo before the constant term..

OpenStudy (anonymous):

Quick question: once I find the square root for the expression inside the sqrt symbol (b - 4ac), do I divide first, or split the equation into two, each with their respective + and - signs?

OpenStudy (campbell_st):

no... since it will be negative.... use i^2 = -1 to help make it a positive value e.g. \[\sqrt{-15} = \sqrt{15i^2} = i \sqrt{15}\]

OpenStudy (campbell_st):

so is the equation the 1 I typed above..?

OpenStudy (anonymous):

Oh, sorry, I think I made a mistake. In the equation, 1200 is added, so it looks like this:\[-3x^4 + 27x^2 + 1200 = 0\]

OpenStudy (campbell_st):

ok... that makes a huge difference divide every term by - 3 and you get \[x^4 - 9x^2 - 400 = 0\] this factors to \[(x^2 - 25)(x^2 + 16) = 0\] now you can go and find all the solutions...

OpenStudy (anonymous):

So the equation might actually come out as x = 4, 5i?

OpenStudy (campbell_st):

well they won't be the solutions.... you need to solve \[x^2 - 25 = 0~~~~~~ and~~~~~~x^2 + 16 = 0\] and there are complex solutions.

OpenStudy (anonymous):

Wait, I think something's wrong with my math. For the quadratic equation, I ended up with:\[x^2 = \frac{ -27 - 123 }{ 6 } , \frac{ -27+123 }{ 6 }\]

OpenStudy (anonymous):

But the answers should be 4i and 5, not 5i and 4?

OpenStudy (anonymous):

Can someone help me figure out what I did wrong?

OpenStudy (campbell_st):

wow.... ok... you need to work through the problem and not work to the solution... the equation is degree 4 so you need to show 4 solutions.... \[x^2 - 25 = 0\] add 25 to both sides of the equation \[x^2 = 25\] solve this and \[x^2 + 16 = 0\] subtract 16 from both sides of the equation \[x^2 = -16\] now solve for x

OpenStudy (anonymous):

x = 4i, and x also equals 5.

OpenStudy (campbell_st):

well they are only 2 of the 4 solutions....

OpenStudy (campbell_st):

\[x^2 = 25~~~~~x = \pm 5\] you can find the other.

OpenStudy (anonymous):

What about the negative forms of the answers, since they would be eliminated after being raised to the 4th and 2nd power?

OpenStudy (campbell_st):

the solutions are \[x = -5, 5 -4i, 4i\] substitute any solution to see if it worked.. you can't have a degree 4 equation with 2 solutions...

OpenStudy (anonymous):

Thank you for your help!